OpenStudy (anonymous):

caleb took a 2 1/2 hour nap after soccer practice on saturday. caleb's younger brother slept 1 3/5 times as longer as caleb. how many hours did his brother sleep

OpenStudy (cwrw238):

can you multiply fractions?

OpenStudy (cwrw238):

2 1/2 * 1 3/5

OpenStudy (cwrw238):

first convert these mixed fractions to improper fractions 2 1/2 = 5/2 (as there are 4 '1/2' s in 2) and 1 3/5 = 8/5

OpenStudy (cwrw238):